OpenStudy (anonymous):

What is the solution set of |x – 6| = 4? 10 10 or –10 –2 or –10 2 or 10

OpenStudy (anonymous):

2 or 10 correct ?

OpenStudy (jiteshmeghwal9):

|a|=a if a>0 & |a|=-a if a<0 therefore, |x-6|=4 has two solutions. -(x-6)=4 & (x-6)=5. solve this two equations.

OpenStudy (jiteshmeghwal9):

what do u gt as the value of x from both equations ?? @rebecca1233

OpenStudy (anonymous):

i got 2 and 10 im just rechecking

OpenStudy (anonymous):

i guess your not going to check it for me but alright bye

OpenStudy (jiteshmeghwal9):

-x+6=4 -x=4-6 -x=-2 x=2 & x-6=4 x=4+6=10

OpenStudy (jiteshmeghwal9):

u r right @rebecca1233 :)